# Three-channel charge Kondo model at high transparency

### v1

@@ Conclusion.
The FRG was used to solve the 3CK model in the quasi-ballistic regime. We confirmed the validity of this approach by recovering, with 2% accuracy, the universal conductance of the 3CK fixed point. The FRG also enabled us to compute the universal energy crossover in the conductance and impurity entropy as functions of frequency or temperature. Furthermore, we extended our study to a model with interacting channels, which connects to the pseudo-gap Kondo model in the low-transparency limit. While electron-electron interactions are difficult to tune in quantum Hall states, they can be enhanced by introducing additional islands, as demonstrated in [68]. Such an experimental setup would enable a test of our predictions in the regime $K\neq 1$.
This case study also provides compelling evidence that the FRG is a powerful method for addressing quantum impurity problems in regimes where conventional techniques are inapplicable.
@@ Acknowledgements.
is $k$ dependent. The expectation value of the field,
@@
$\partial_{t}U_{k}({\bm{\phi}})=\frac{1}{2\beta}\sum_{\omega_{n}}e^{-\omega_{n}^{2}/D^{2}}\int d^{2}\phi\,{\rm tr}[\partial_{t}R_{k}(i\omega_{n})G_{k}(i\omega_{n},{\bm{\phi}})],$ (S36)
where $\int d^{2}\phi$ is the integral over a unit cell of the triangular lattice spanned by the vectors ${\bm{a}}_{1}$ and ${\bm{a}}_{2}$ (defined after (S17)). The trace is taken with respect to the indices of the $2\times 2$ matrices $R_{k}(i\omega_{n})$ and $G_{k}(i\omega_{n},{\bm{\phi}})$. The propagator $G_{k}(i\omega_{n},{\bm{\phi}})$ is obtained by inverting $\Gamma^{(2)}_{k}(i\omega_{n},{\bm{\phi}})+R_{k}(i\omega_{n})$, where $\Gamma^{(2)}_{k}(i\omega_{n},{\bm{\phi}})$ is the two-point vertex computed in a constant field $\bm{\phi}$}. The factor $e^{-\omega_{n}^{2}/D^{2}}$ in (S36) implements the UV energy cutoff. In practice, the flow equation (S36) is solved using the dimensionless quantities $\tilde{\omega}_{n}=\omega_{n}/k$ and $\tilde{r}_{k}^{(n,m)}=r_{k}^{(n,m)}/k$.
@@
The LPA allows us to compute the free energy $F=U_{k=0}({\bm{\phi}}=0)$ and the entropy $S=-\partial F/\partial T$. However, it does not allow for the calculation of the conductance, $G(\omega)$, which requires knowledge of the full frequency dependence of the propagator $G_{k=0,ij}(i\omega_{n})=\langle\varphi_{i}(i\omega_{n})\varphi_{j}(-i\omega_{n})\rangle$; see Eq. (S26). A derivative expansion of the effective action beyond the LPA is subject to the same limitations. The propagator $G_{k=0}$, or its inverse the two-point vertex $\Gamma^{(2)}_{k=0}$, can be computed within the BMW approximation [18, 19, 20]. The latter allows one to close the infinite hierarchy of equations satisfied by the vertices $\Gamma^{(n)}_{k}$. In its simplest nontrivial implementation, the BMW approximation yields closed equations for the effective potential $U_{k}({\bm{\phi}})$ and the two-point vertex $\Gamma^{(2)}_{k}(i\omega_{n},{\bm{\phi}})$ evaluated in a constant field. The latter can be written as
@@ Appendix A Appendix A: Numerical integration of the flow equations
$+I_{2,k}^{i_{1}i_{2}i_{3}i_{4}}(i\omega)\bigl{[}|\omega|\partial_{\phi_{i_{2}}}X_{k,ii_{3}}\partial_{\phi_{i_{1}}}X_{k,ji_{4}}+\partial_{\phi_{i_{2}}}X_{k,ii_{3}}U^{(3)}_{k,i_{1}ji_{4}}+\partial_{\phi_{i_{1}}}X_{k,ji_{4}}U^{(3)}_{k,i_{2}ii_{3}}\bigr{]}\biggr{\}}$. (S50)

### v2

@@ Conclusion.
In summary, we have solved the three-channel charge Kondo model in the quasi-ballistic regime using the functional renormalization group. Our results recover the universal zero-frequency conductance of the three-channel Kondo fixed point with excellent quantitative accuracy and provide the full universal crossover of the conductance and impurity entropy as functions of temperature or frequency. Extending the analysis to interacting channels, we uncovered a continuous line of fixed points that connects smoothly to pseudo-gap Kondo physics in the low-transparency limit. While tuning electron–electron interactions in quantum Hall systems is challenging, recent experiments with multiple islands demonstrate viable routes to enhance them, opening possibilities to test our predictions for $K\neq 1$ [67, 68]. More broadly, this work establishes FRG as a powerful nonperturbative approach to quantum impurity problems in regimes inaccessible to conventional methods, with direct implications for future experiments in mesoscopic devices.
@@ Acknowledgements.
is $k$-dependent. The expectation value of the field,
@@
$\partial_{t}U_{k}({\bm{\phi}})=\frac{1}{2\beta}\sum_{\omega_{n}}e^{-\omega_{n}^{2}/D^{2}}\,{\rm Tr}[\partial_{t}R_{k}(i\omega_{n})G_{k}(i\omega_{n},{\bm{\phi}})],$ (S36)
The trace is taken with respect to the indices of the $2\times 2$ matrices $R_{k}(i\omega_{n})$ and $G_{k}(i\omega_{n},{\bm{\phi}})$. The propagator $G_{k}(i\omega_{n},{\bm{\phi}})$ is obtained by inverting $\Gamma^{(2)}_{k}(i\omega_{n},{\bm{\phi}})+R_{k}(i\omega_{n})$, where $\Gamma^{(2)}_{k}(i\omega_{n},{\bm{\phi}})$ is the two-point vertex computed in a constant field $\bm{\phi}$}. The factor $e^{-\omega_{n}^{2}/D^{2}}$ in (S36) implements the UV energy cutoff. In practice, the flow equation (S36) is solved using the dimensionless quantities $\tilde{\omega}_{n}=\omega_{n}/k$ and $\tilde{r}_{k}^{(n,m)}=r_{k}^{(n,m)}/k$.
@@
The LPA allows us to compute the free energy $F=U_{k=0}({\bm{\phi}}=0)$ and the entropy $S=-\partial F/\partial T$. However, it does not allow for the calculation of the conductance, $G(\omega)$, which requires knowledge of the full frequency dependence of the propagator $G_{k=0,ij}(i\omega_{n})=\langle\varphi_{i}(i\omega_{n})\varphi_{j}(-i\omega_{n})\rangle$; see Eq. (S26). A derivative expansion of the effective action beyond the LPA is subject to the same limitations. The propagator $G_{k=0}$, or its inverse the two-point vertex $\Gamma^{(2)}_{k=0}$, can be computed within the BMW approximation [18, 19, 20, 21]. The latter allows one to close the infinite hierarchy of equations satisfied by the vertices $\Gamma^{(n)}_{k}$. In its simplest nontrivial implementation, the BMW approximation yields closed equations for the effective potential $U_{k}({\bm{\phi}})$ and the two-point vertex $\Gamma^{(2)}_{k}(i\omega_{n},{\bm{\phi}})$ evaluated in a constant field. The latter can be written as
@@ Appendix A Appendix A: Numerical integration of the flow equations
[…]